Determine the amount of heat that must be removed from 1 kg of water at 40°C to convert it to ice at -5°C.

(Obtain any relevant data that you need from the internet. Cite the source of that data in your answer)


1
Expert's answer
2022-01-18T09:35:16-0500

To cool the water down from 40 to 0°C:


"Q_1=c_1m(t_1-t_0)=4200·1·(40-0)=168000\text{ J}."


To freeze the 1 kg of water at 0°C [2]:


"Q_2=m\lambda=1·334000=334000\text{ J}."


To cool down the 1 kg of ice from 0 to -5° [3]:


"Q_3=c_2m(t_0-t_2)=2100·1·(0-(-5))=10500\text{ J}."


The total amount of heat:


"Q=Q_1+Q_2+Q_3=512.5\text{ kJ}."


Sources:

[1] https://www.bbc.co.uk/bitesize/guides/zchgqhv/revision/3

[2] https://www.britannica.com/science/latent-heat

[3] https://gchem.cm.utexas.edu/data/section2.php?target=heat-capacities.php